OpenStudy (anonymous):

-w/2=-15 solve for w giving steps please

OpenStudy (across):

\[-\frac{w}{2}=-15\]\[-\frac{w}{2}\cdot2=-15\cdot2\]\[-\frac{w}{\not{2}}\cdot\not{2}=-30\]\[-w=-30\]\[-w\cdot-1=-30\cdot-1\]\[w=30\]